OpenStudy (anonymous):

Consider the reaction. N2(g) + 3H2(g) <-> 2NH3(g) At equilibrium, the concentrations of the different species are as follows. [NH3] = 0.105 M [N2] = 1.1 M [H2] = 1.50 M What is the equilibrium constant for the reaction at this temperature? 0.0030 0.030 34 340

OpenStudy (anonymous):

The way to calculate the equilibrium constant is you multiply the concentrations of product over the products of reactants, with their coefficients as exponents. That sounds weird, but look it's not so bad! \[K = \frac{[NH_3]^2}{[N_2][H_2]^3}\]

OpenStudy (rushwr):



OpenStudy (photon336):

Also another thing to point out

OpenStudy (photon336):

they said EQUILIBRIUM CONCENTRATIONS. (sorry for the caps) which means that your reaction has already reached equilibrium, so as @woodard and @rushwr said you would set up your equilibrium constant, and plug in the numbers for the concentrations. but if the reaction did not reach equilibrium, and those are NOT the equilibrium concentrations, then you need to do an ice table, to show how the concentrations change over time, but technically you would need the equilibrium constant to do this Keq. say if you had that reaction below: \[AB(g)--> C(g) + D(g)\] \[K = \frac{ [C][[D] }{ [A][B]}\] YOU would raise the concentrations to the power of the coefficient in which it appears in your chemical formula. for A,B,C, and D. the coefficient is 1.